What will be the result if a student connect voltmeter in a series and a ammeter in parallel in a circuit by mistake?

Grade:Upto college level

2 Answers

Abhi Avi
7 years ago
By the presence of voltmeter in series with resistance R in a circuit due to voltmeter`s high resistance the current I(=V/R) will get decreases to I`1(=V/R) and as I`1 will divide in resistance and ammeter the current through ammeter I`2 will ever lesser than I`1. Also the P.D will divide across the resistance and voltmeter, so the P.D across V`2 will lesser than V. Since I`2 and V`2 both are lesser than I and V respectively, none of the instruments will be damage
Shamreen Jahan
6 years ago
voltmeter have high resistane so when it is connected in series it decreases the current going to the circuit.while ammeter has low resistnce,when it is connected in parallel current would bypass the circuit .all go through ammeter and ammeter will burn out.